Antonio is buiklding a bookshelf. He has a piece of wood that measures 90 centimeters in length. The directions require that this piece of wood be cut into two pieces, one piece 20 centimeters longer than the other. He uses the equation below to determine where to make the cut.
X+(X+20)=90
The shorter piece of wood must then be cut in half. Use the equation above to help you determine the length of the shorter piece of wood.
Show you work.

The answer I have is:
25+(45+20)=90
25 centimeters for the shortest piece of wood
The answer I have is wrong, not sure what to do. Please help....
Thank You
7th grader in NY

Answer by doctor_who(15)

See if this helps ..
Let the length of the shorter piece of wood be x centimetres in length.
The longer piece is 20cm longer, so it is (x+20) centimentres long.
.. and the total length of the two pieces put together is 90cm.
so : x + (x+20) = 90 [as you rightly say]
In this case, we can dispense with the brackets
x + x + 20 = 90
Subtract 20 from each side :
x + x = 70
Which is another way of saying :
2x = 70
Now halve both sides :
x = 35 cm (ANSWER)
If this piece of wood is then cut in half again, then each piece will be 17.5 cm long (ANSWER)
(if I understand the question correctly)
Hope this helps!